How many ways can 3 outfielders and 4 infielders be chosen from 5 outfielders and 7 infielders.
Answer by stanbon(48568)

How many ways can 3 outfielders and 4 infielders be chosen from 5 outfielders and 7 infielders.
--------------------------------
Choose 3 outfielder: 5C3 = (5*4)/(1*2) = 10
Choose 4 infielders: 7C4 = (7*6*5)/(1*2*3) = 35
--------------------------------
Total number of ways to pick outfielders and infielders: 10*35 = 350
=======================================
Cheers,
Stan H.